A bus takes 8 houre to cover a distance of 320 km/h what is the average speed of the bus

Grade:9

8 Answers

Nitesh
6 years ago
Answer would be 40 km/hour if distance was written 320 km and total time is 8 hour.Then average velocity=total distance travelled/total time taken=320/8=40km/hour.So the answer is 40 km/hour.Please correct the question first.Write 320 km instead of 30 km/ hour.
Akhilesh shukla
6 years ago
For ans time given should be converted into hr and distance in m/sec.then time becomes 480seb and distance becomes 88.8m/sec then avg speed=total dist/time,therefore 88.8/480=0.185m/sec.is the ans.
Nitesh
6 years ago
We can convert these quantity into another another unit at any time either during answer or after getting answer.So the answer also would be 40 km/hr or 100/9 m/ sec.My answer is correct.
Kuldip Pawar
6 years ago
Avg.Velocity=total distance/total time.Therefore=--->320Km/8hours.Is equal to=--->40Km/hours.We can convert this answer in m/s according to option
Milind
6 years ago
I think there is problems in your question. Form which book you given this question. Please recheck the question 
Abid
5 years ago
first of all we convert the units of these given quantities in standard unit(S.I) then by formula we solve it.
data give ;t= 8hours=8×60mins×60secs=28800secs
  distance=320km/h=320×1000m=320000m
formula;  Vaverage=distance traveled÷time taken
                             =320000m÷28800sec
                             =11.11m/s
yashveer
5 years ago
As we know that ,
average speed=total distance/total time
                        =320 km / 8 hr
                        =40 km/hr
note:please correct question at place of “ / “ write per because it seems that you have written km/hr (velocity)
Durwa Thakur
5 years ago
Average speed= 320km/8h.
So, avg speed= 40km/h.
Hope this helps.
Regards.
Btw, in the question you've written 320km/h but it should be 320 km. I've assumed it's 320 km.